Solve this Aggregate Planning Problem by minimizing the cost of matching the capacity various options in various periods to tAll Periods Total Regular Time Cost 52800 Total Overtime Cost 6200 Total Subcontracting Cost Total Cost for All 59000 cheek the answer plz
Answer #1
Period 1 Period 2 Period 3 Excess Capacity Row Total
Beginning inventory 0 0 0 0 0
Period 1 reg time 700 100 200 0 1000
Period 1 overtime 0 0 0 100 100
Period 1 subcontracting 0 0 0 200 200
Period 2 reg time 0 800 0 0 800
Period 2 overtime 0 100 0 0 100
Period 2 subcontracting 0 0 0 200 200
Period 3 reg time 0 0 800 0 800
Period 3 overtime 0 0 100 0 100
Period 3 subcontracting 0 0 0 100 100
Column Total 700 1000 1100 600 3400
Total Cost 59000

Regular time cost = 700*20+100*22+200*24+800*20+800*20=53000

Overtime cost = (100+100)*30= 6000

Total subcontracting cost = 0

Please rate me

Thanks